# Review on the decoding algorithms for surface codes

### v4

@@ 2.2 The decoding problem
    where $P$ refers to the probability distribution function of the errors. Due to the fact that degeneracy is ignored by this type of decoding, it is also referred as non-degenerate decoding.
  • Degenerate quantum maximum likelihood decoding (DQMLD): due to the existence of degenerate errors that form cosets of errors that affect the coded state in a similar manner, it is possible that a the probability of occurrence of the coset containing the most probable error sequence (in the sense of QMLD) is smaller than other coset allowed by the measured syndrome. Thus, the QMLD decoder will be suboptimal as it ignores the degenerate structure of stabilizer codes. Therefore, DQMLD decoding can be described mathematically as [54, 55]:
@@
Therefore, the optimal decoding rule for stabilizers is the DQMLD. However, it was proven that QMLD falls into the NP-complete complexity class (similar to the classical decoding problem), while DQMLD belongs to the #\#P-complete class [55, 80, 81, 82]. The latter is computationally much harder than the other, implying that the optimal decoding rule may pose serious issues for the fast decoding needed in quantum error correction [55]. Therefore, even if the optimal rule for decoding and, thus, code performance is obtained by using DQMLD, non-degenerate decoding is important and widely used as it is less expensive in terms of computational complexity.
3 The surface Code
@@
Figure 1: Distance-7 planar code. Data qubits are represented as grey circles, $X$ and $Z$-checks are represented as green and orange circles respectively.
@@
Surface codes are a family of quantum error correcting codes in which the information of logical qubits is mapped to a set of physical qubits, commonly named as data qubits, which are displayed in a lattice array. Moreover, the measurement qubits that are used to measure the checks are also displayed within the lattice. Alexei Kitaev first proposed the concept of surface codes in his prominent work [40], where the qubits were displayed in a torus-shaped lattice. This toric code has periodic boundary conditions and it is able to protect two logical qubits. Nevertheless, such toric displacement of the qubits makes the hardware implementation and logical qubit connectivity complicated [83] since many experimental implementations such as superconducting hardware, for example, may require the system to be placed in a two-dimensional lattice. Thus, the so called planar code encoding a single logical qubit is obtained by stripping the periodic boundary conditions from the toric code [32, 84]. Specifically, it will be a $d^{2}+(d-1)^{2},1,d]]$ QECC. Furthermore, the number of data and measurement qubits used for protecting the single logical qubit can be lowered down, i.e. the rate of the code can be increased [^8] The rate of a quantum error correction code is defined as the ratio between the number of logical qubits and the number of physical qubits, i.e. $R_{\mathrm{Q}}=k/n$, by considering a specific set of data qubits and checks within the planar code. The obtained code is usually known as the rotated planar code, which will be a $d^{2},1,d]] [$62]. FIG.1 shows a distance-7 planar code where the dashed lines indicate the subset of qubits that form the rotated planar code with the same distance. In this this tutorial we will consider the rotated planar code in a square lattice with Calderbank-Shor-Steane (CSS) structure  [^9] CSS codes refer to stabilizer codes admitting a set of generators that are either $X$ or $Z$-check operators. This means that the check operators will consist of either tensor products of identities with $X$ or with $Z$ operators exclusively [85, 86]. due to its practicality and relevance at the time of writing. Note that this code is the one that has been recently implemented experimentally by Wallraff’s group at ETH Zurich [63] and by the Google Quantum AI team [64]. Nevertheless, the decoding methods discussed in this tutorial apply for all surface codes including the tailored versions proposed through the past years [41, 42, 87, 88] or the different lattices considered [62, 43, 89].
@@ 3.1 Stabilizer and check structure
Due to the structure of their stabilizers, CSS quantum surface codes have two types of checks: $X$-checks and $Z$-checks. The former detect $X$-errors, while the latter detect $Z$-errors. FIG. 2 portrays the structure of the check operators in a distance-5 CSS rotated planar code. As it can be seen in the figure, this structure fulfills the condition that the stabilizer generators form an abelian group [39]. The locality of the check operations ensures that checks that are far apart commute with each other, while adjacent checks commute either because they are of the same type and, thus, apply the same operators to their adjacent data qubits; or because they anti-commute for two data qubits at the same time, making the whole operators to commute. For example,
Figure 2: Distance-5 rotated planar code. The stabilizing operators that the checks yield over their adjacent data qubits are denoted through light green and orange circles.
@@ 3.2 Types of errors and code threshold
In FIG.4 we show some examples of errors that may arise in the rotated planar code. In the upper right section of the code, there are three isolated Pauli errors, namely an $X$, a $Y$ and a $Z$-error, which cause the adjacent susceptible checks to exhibit non-trivial syndrome elements upon measurement. Note that the $Y$-error triggers both the $X$ and $Z$-checks that are adjacent to such data qubit. This happens because $Y$-errors are a combination of $X$ and $Z$-errors, neglecting the global phase, as seen in eq.(3). These isolated Pauli errors will be detected by the code, and using the syndrome information, the decoder will try to estimate which are those errors. Note that these errors have weight one, and since this specific rotated planar code has distance-$5$, those errors can be corrected.
@@
In addition, two other errors forming a vertical and horizontal chain along the left and bottom boundaries are represented in FIG. 4. As seen in the figure, those error chains are not detected by the code since they do not trigger any of the surrounding measurement qubits. This is because each susceptible check is connected to two of the Pauli operators, i.e. it refers to the previously described case where there is an even number of operators acting on each of the checks. This error chains act non-trivially on the codestate without being part of the stabilizer group while presenting a trivial syndrome, i.e. they belong to the normalizer of the code. This type of errors receive the name of logical errors [32]. Specifically, the left error chain is a logical $Z$-error, $\mathrm{Z}_{L}$, the bottom error chain is a logical $X$-error, $\mathrm{X}_{L}$, and the combination of them two is a logical $Y$-error, $\mathrm{Y}_{L}$. Notice that the anti-commutation relation $\mathrm{X}_{L}\mathrm{Z}_{L}=-\mathrm{Z}_{L}\mathrm{X}_{L}$ is preserved through the bottom left data qubit. Note that these are just two examples of logical errors, there are other possible logical operators. Moreover, if a logical error is applied, the resulting state will still be within the codespace, so eq.(4) will still be preserved for the new sate and it will remain invariant upon the application of stabilizer operators. This can be reflected in the bottom right of FIG.4, where the application of the bottom left $Z$-check modifies the shape of the $\mathrm{X}_{L}$ operator while still commuting with all the checks.
@@
Figure 4: The figure shows a CSS rotated planar code, with physical Pauli operators represented by red circles. Check measurements resulting in non-trivial syndrome elements are marked with an exclamation mark. The thick red line represent the shape of two logical operators, while the pink dashed line shows the changed path of one of them after interacting with a stabilizer.
@@
When protecting against quantum noise, experiencing a logical error [^10] Note that here we refer to the unwanted event that the codestate is altered by a logical error due to the noise. However, whenever the logical state is wanted to be manipulated with the scope of computation, the way of applying logical Pauli gates to the logical qubit is by means of these logical operators on the physical qubits. is fatal since those alter the information stored in the code without any indication from the checks and, thus, cannot be detected nor corrected. One way of mitigating the impact of logical errors is to increase the size of the surface code. By doing so, the code distance increases and, therefore, the minimum number of Pauli operators needed to form a logical error will also be higher, making such event to be more improbable. This comes from the fact that logical errors belong to the normalizer of the code and the distance is defined as the minimum weight of those Pauli operators. Although it may seem intuitive that a larger surface code would perform better, this is not always the case. The momentous result in quantum computing named the threshold theorem states that the performance of a QECC improves as its distance increases, provided that the physical error rate of the data qubits is below a certain value [29, 40, 90, 91]. This value is known as the probability threshold ($p_{th}$). As long as the physical error rate is below $p_{th}$, increasing the size of the surface code will lead to a better code performance.
@@ 4 Noise models
The main obstacle to construct quantum computers is the proneness of quantum information to suffer from errors. There are many error sources that corrupt quantum information while being processed such as state preparation and measurement (SPAM) errors or errors introduced by imperfect implementations of quantum gates, for example. Many of the errors occur due to the fact that the technology used for manipulating qubits is imperfect [32, 33]. However, qubits do also suffer from errors due to their undesired and unavoidable, in principle, interaction with their surrounding environment. This last source of errors is named as environmental decoherence, and corrupts quantum information even if the quantum system is left to evolve freely [20, 34, 30, 35]. Thus, decoherence poses a fundamental problem to the field of quantum information processing since its existence does not depend on imperfect implementations of qubit manipulations or measurements, which we may deem as engineering problems [^11] It may seem unfair to state that decoherence is not related to engineering since the way qubits are constructed fundamentally determines how fast a qubit will decohere. However, even in the case that very long decoherence times are obtained, it would not be possible to apply an arbitrarily large amount of perfect quantum gates, since at some point the quantum information would be corrupted.. Therefore, quantum error correction will be necessary if we want to run arbitrarily large quantum algorithms with enough precision so that the obtained results are reliable, even if perfect quantum gates, measurements and state preparations are available.
@@ 4.1 Decoherence
In general, decoherence comprises several physical processes that describe the qubit environment interaction, and the nature of those processes depends on the qubit technology, i.e. superconducting qubits, ion traps or NV centers, for example. However, most of those physical interactions can be grouped into three main decoherence mechanisms [^12] Here we are considering noise sources that operate on the computational subspace of the qubit, i.e. transitions to other possible levels are neglected by now. This will be described later on. since their operational effect on the two-level coherent system that is the qubit is the same [30, 92]:
@@
There are many ways in which this set of physical interactions can be mathematically described such as the Gorini-Kossakowski-Lindblad-Sudarshan (GKLS) master equation [94, 95, 96] or the quantum channel formulation [30, 97]. In the context of QECC, quantum channels are used to describe noisy evolution. In general, quantum channels are linear, completely-positive, trace-preserving (CPTP) maps between spaces of operators. As a consequence of those properties, quantum channels fulfill the Choi-Kraus theorem, implying that the application of such maps on a density operator $\rho$ can be written as the following decomposition [20, 30, 97]:
@@
The generalized amplitude and phase damping channel (GAPD), $\mathcal{N}_{\mathrm{GAPD}}$, describes the evolution of a quantum state when decoherence arises from the three qubit-to-environment interactions presented before [30, 35, 93, 98, 99]. Such channel consists of a generalized amplitude damping channel (GAD) describing the thermal and relaxation interaction [93, 98, 99] and of a pure dephasing channel (PD) [30, 35]. The action of the GAD is described by the damping parameter, $\gamma$, and the probability that the ground state is excited by finite temperature, $N$. The damping parameter relates to the relaxation time of the qubit as, $\gamma(t)=1-e^{-(2n_{\mathrm{th}}+1)t/T_{1}}$, with $t$ being the evolution time, while $n_{\mathrm{th}}$ and $N(n_{\mathrm{th}})$ depend on the temperature and the energy gap of the system [93, 98, 99]. Whenever the thermal excitation is considered negligible, $n_{\mathrm{th}}\approx 0$ and $N(n_{\mathrm{th}})=0$; and such channel reduces to an amplitude damping channel (AD). The pure dephasing channel (PD) is described by the so called scattering probability, $\lambda$, which relates to the pure dephasing time as, $\lambda(t)=1-e^{-2t/T_{\phi}}$, with $t$ the evolution time again [35, 78]. In this sense, the GAPD channel is defined by those parameters.

### v5

@@ 2.2 The decoding problem
    where $P$ refers to the probability distribution function of the errors. Due to the fact that degeneracy is ignored by this type of decoding, it is also referred as non-degenerate decoding. For an independent, identically distributed noise model, this decoding rule results in a minimum-weight decoding rule, i.e. looking for the error with minimum weight that satisfies the measured syndrome.
  • Degenerate quantum maximum likelihood decoding [^9] Note that the literature sometimes names this decoding rule as “maximum likelihood decoding” [60]. However, both rules are actually maximum likelihood estimation problems. Their difference is the likelihood function to maximize. Therefore, we stick to the terminology in [42]. (DQMLD): due to the existence of degenerate errors that form cosets of errors that affect the coded state in a similar manner, it is possible that a the probability of occurrence of the coset containing the most probable error sequence (in the sense of QMLD) is smaller than other coset allowed by the measured syndrome. Thus, the QMLD decoder will be suboptimal as it ignores the degenerate structure of stabilizer codes. Therefore, DQMLD decoding can be described mathematically as [41, 42]:
@@
Therefore, the optimal decoding rule for stabilizers is the DQMLD. However, it was proven that QMLD falls into the NP-complete complexity class (similar to the classical decoding problem), while DQMLD belongs to the #\#P-complete class [42]. The latter is computationally much harder than the other, implying that the optimal decoding rule may pose serious issues for the fast decoding needed in quantum error correction [42]. Therefore, even if the optimal rule for decoding and, thus, code performance is obtained by using DQMLD, non-degenerate decoding is important and widely used as it is less expensive in terms of computational complexity.
3 The surface Codes
@@
Figure 1: Distance-7 planar code. Data qubits are represented as white circles, $X$ and $Z$-checks are represented as green and orange circles respectively.
@@
Surface codes are a family of quantum error correcting codes in which the information of logical qubits is mapped to a set of physical qubits, commonly named as data qubits, which are displayed in a lattice array. Moreover, the measurement qubits that are used to measure the checks are also displayed within the lattice. Alexei Kitaev first proposed the concept of surface codes in his prominent work [31], where the qubits were displayed in a torus-shaped lattice. This toric code has periodic boundary conditions and it is able to protect two logical qubits. Nevertheless, such toric displacement of the qubits makes the hardware implementation and logical qubit connectivity complicated [61] since many experimental implementations such as superconducting hardware, for example, may require the system to be placed in a two-dimensional lattice. Thus, the so called planar code encoding a single logical qubit is obtained by stripping the periodic boundary conditions from the toric code [25, 62]. Specifically, it will be a $d^{2}+(d-1)^{2},1,d]]$ QECC. Furthermore, the number of data and measurement qubits used for protecting the single logical qubit can be lowered down, i.e. the rate of the code can be increased [^10] The rate of a quantum error correction code is defined as the ratio between the number of logical qubits and the number of physical qubits, i.e. $R_{\mathrm{Q}}=k/n$, by considering a specific set of data qubits and checks within the planar code. The obtained code is usually known as the rotated planar code, which will be a $d^{2},1,d]] [$49]. FIG. 1 shows a distance-7 planar code where the dashed lines indicate the subset of qubits that form the rotated planar code with the same distance. In this tutorial we will consider the rotated planar code in a square lattice with Calderbank-Shor-Steane (CSS) structure  [^11] CSS codes refer to stabilizer codes admitting a set of generators that are either $X$ or $Z$-check operators. This means that the check operators will consist of either tensor products of identities with $X$ or with $Z$ operators exclusively [63, 64]. due to its practicality and relevance at the time of writing. Note that this code is the one that has been recently implemented experimentally by Wallraff’s group at ETH Zurich [50] and by the Google Quantum AI team [51]. Nevertheless, the decoding methods discussed in this tutorial apply for all surface codes including the tailored versions proposed through the past years [32, 33, 65, 66] or the different lattices considered [49, 34, 67]. From here through the end of the review, we will use the terms rotated planar code and rotated code interchangeably so as not to sound repetitive.
@@ 3.1 Stabilizer and check structure
Due to the structure of their stabilizers, CSS quantum surface codes have two types of checks: $X$-checks and $Z$-checks. The former detect $X$-errors, while the latter detect $Z$-errors. FIG. 2 portrays the structure of the check operators in a distance-3 CSS rotated planar code. As it can be seen in the figure, this structure fulfills the condition that the stabilizer generators form an abelian group [30]. This can be seen by explicitly studying the stabilizer generators:
$s_{1}$ $=Z_{2}Z_{3},$ (9)
$s_{2}$ $=Z_{1}Z_{2}Z_{4}Z_{5},$
$s_{3}$ $=Z_{5}Z_{6}Z_{8}Z_{9},$
$s_{4}$ $=Z_{7}Z_{8},$
$s_{5}$ $=X_{1}X_{4},$
$s_{6}$ $=X_{4}X_{5}X_{7}X_{8},$
$s_{7}$ $=X_{2}X_{3}X_{5}X_{6},$
$s_{8}$ $=X_{6}X_{9},$
where $s_{i}$ denotes the $i$th stabilizer generator or check. The locality of the check operations ensures that checks that are far apart commute with each other, while adjacent checks commute either because they are of the same type and, thus, apply the same operators to their adjacent data qubits; or because they anti-commute for two data qubits at the same time, making the whole operators to commute. For example,
Figure 2: Distance-3 rotated planar code. The stabilizing operators that the checks yield over their adjacent data qubits are denoted through green and orange squares, which act on their nearest data qubits.
@@
As explained before, the data qubits of surface codes may undergo a Pauli error. The syndrome of the associated error is measuring the check operators, which corresponds to the quantum circuits shown in FIG. 3. The top circuit represents an $X$-check and the bottom circuit a $Z$-check. As seen in such figure, if an odd number of adjacent data qubits suffer from an $X$ or $Z$-error, the measurement of the respective $X$ or $Z$-checks will be triggered. However, as seen in the top image from FIG. 3, in the event of an even number of errors, those will cancel, due to their unitary nature, and no error will be detected by the check operator measurement.
@@ 3.2 Types of errors and code threshold
In FIG. 4 we show different types of errors that may arise in the rotated planar code. Beginning with the top figure, in the upper right section of the code, there are three isolated Pauli errors, namely an $X$, a $Y$ and a $Z$-error, which cause the adjacent susceptible checks to exhibit non-trivial syndrome elements upon measurement. Note that the $Y$-error triggers both the $X$ and $Z$-checks that are adjacent to such data qubit. This happens because $Y$-errors are a combination of $X$ and $Z$-errors, neglecting the global phase, as seen in eq.(3). These isolated Pauli errors will be detected by the code, and using the syndrome information, the decoder will attempt to estimate them.
@@
In addition, other errors forming vertical and horizontal chains along the boundaries are represented in all figures of FIG. 4 and denoted with bold red lines. As seen in the figure, those error chains are not detected by the code since they do not trigger any of the surrounding measurement qubits. This is because each susceptible check is connected to two of the Pauli operators, i.e. it refers to the previously described case where there is an even number of operators acting on each of the checks. These error chains act non-trivially on the codestate without being part of the stabilizer group while presenting a trivial syndrome, i.e. they belong to the normalizer of the code. This type of errors receive the name of logical errors [25]. Specifically, in FIG. 4, both the vertical and horizontal chains from the top figure, the horizontal chain from the middle figure and the vertical chain from the bottom figure constitute the $Y_{L},X_{L}$ and $Z_{L}$ operators respectively. Notice that, in the top figure, the anti-commutation relation $\mathrm{X}_{L}\mathrm{Z}_{L}=-\mathrm{Z}_{L}\mathrm{X}_{L}$ is preserved through the bottom left data qubit. Moreover, these are just three examples of the physical Pauli operators producing the logical errors. If a logical operator is applied, the resulting state will still be within the codespace, so eq.(4) will still be preserved for the new sate and it will remain invariant upon the application of stabilizer operators. This can be reflected in all three figures within FIG. 4, where the application of three arbitrary stabilizers (one for each figure), produces another chain of Pauli operators spanning from one boundary to the other, commuting with the stabilizer set. The new operator, which is denoted by a dashed line and would act non-trivially on the data qubits it goes over, is different in terms of physical Pauli operators, but acts equally on the codespace.
@@
Figure 4: The figure illustrates three rotated codes undergoing a logical $Y_{L},X_{L}$ and $Z_{L}$ operator from top to bottom. Physical non-trivial Pauli operators represented by red circles. Check measurements resulting in non-trivial syndrome elements are marked with an exclamation mark. The red thick line represent the shape of logical operators, while the red dashed line shows the changed path of a logical operator after interacting with a stabilizer the physical qubit operators of which are denoted as squares.
@@
When protecting against quantum noise, experiencing a logical error [^12] Note that here we refer to the unwanted event that the codestate is altered by a logical error due to the noise. However, whenever the logical state is wanted to be manipulated with the scope of computation, the way of applying logical Pauli gates to the logical qubit is by means of these logical operators on the physical qubits. is fatal since those alter the information stored in the code without any indication from the checks and, thus, cannot be detected nor corrected. One way of mitigating the impact of logical errors is to increase the size of the surface code. By doing so, the code distance increases and, therefore, the minimum number of Pauli operators needed to form a logical error will also be higher, making such event to be more improbable. This comes from the fact that logical errors belong to the normalizer of the code and the distance is defined as the minimum weight of those Pauli operators. Although it may seem intuitive that a larger surface code would perform better, this is not always the case. The momentous result in quantum computing named the threshold theorem states that, by means of quantum error correction, the logical error rate experienced by a quantum computer can be suppressed to an arbitrarily low level provided that the physical error rate of the data qubits is below a certain value [24, 31, 68, 69]. This closely relates with the concept of code threshold ($p_{th}$), which refers to the maximum physical error rate for which increasing the distance of a code in consideration results in a better performance for a certain error model. Note that not all QECCs present a code threshold, e.g. the Bacon-Shor code [70], implying that increasing distance does not always result in lowering the logical error rate. Note that not having threshold is equivalent to saying that the threshold is zero. The considered rotated planar surface code presents non-trivial code thresholds for the noise models considered in theoretical QEC. Thus, as long as the physical error rate is below $p_{th}$, increasing the size of such code will lead to a better performance.
@@ 4 Noise models
The main obstacle to construct quantum computers is the proneness of quantum information to suffer from errors. There are many error sources that corrupt quantum information while being processed such as state preparation and measurement (SPAM) errors or errors introduced by imperfect implementations of quantum gates, for example. Many of the errors occur due to the fact that the technology used for manipulating qubits is imperfect [25, 26]. However, qubits do also suffer from errors due to their undesired interaction with their surrounding environment. This last source of errors is named as environmental decoherence, and corrupts quantum information even if the quantum system is left to evolve freely [15, 28, 27, 29]. Thus, decoherence poses a fundamental problem to the field of quantum information processing since its existence does not depend on imperfect implementations of qubit manipulations or measurements, which we may deem as engineering problems [^13] It may seem unfair to state that decoherence is not related to engineering since the way qubits are constructed fundamentally determines how fast a qubit will decohere. However, even in the case that very long decoherence times are obtained, it would not be possible to apply an arbitrarily large amount of perfect quantum gates, since at some point the quantum information would be corrupted.. Therefore, quantum error correction will be necessary if we want to run arbitrarily large quantum algorithms with enough precision. It is noteworthy to state that there are approaches to construct qubits that are naturally protected from decoherence by means of topological order or physical symmetries [43]. These methods are usually referred to as passive quantum error correction. However, in this review we are focusing on active error correction, in the sense that errors are corrected by actively detecting and correcting them.
The aim of this section is to describe the usual noise sources considered in theoretical quantum error correction as well as the motivation of certain error models, e.g. biased noise. In this sense, the usual error models are the code capacity noise, phenomenological noise and circuit-level noise. However, it is important to state that many important error sources are usually neglected by this models such as coherent errors, non-Markovian noise and correlated errors arising from crosstalk or cosmic rays (in superconducting qubits) [45, 51, 71, 72].
@@ 4.1 Decoherence
In general, decoherence comprises several physical processes that describe the qubit environment interaction, and the nature of those processes depends on the qubit technology [^14] We consider conventional qubit-based quantum computation, i.e. discrete two-level coherent quantum systems., i.e. superconducting qubits, ion traps or NV centers, for example. However, most of those physical interactions can be grouped into three main decoherence mechanisms [^15] Here we are considering noise sources that operate on the computational subspace of the qubit, i.e. transitions to other possible levels are neglected by now. This will be described later on. since their operational effect on the two-level coherent system that is the qubit is the same [27, 73]:
@@
For Markovian [^17] Markovian noise assumes that the system-environment interaction is memoryless [43]. noise, there are many ways in which this set of physical interactions can be mathematically described such as the Gorini-Kossakowski-Lindblad-Sudarshan master equation [75, 76] or the quantum channel formulation [27, 77]. In the context of QECC, quantum channels are used to describe noisy evolution. In general, quantum channels are linear, completely-positive, trace-preserving maps between spaces of operators. As a consequence of those properties, quantum channels fulfill the Choi-Kraus theorem, implying that the application of such maps on a density operator $\rho$ can be written as the following decomposition [15, 27, 77]:
@@
The generalized amplitude and phase damping channel describes the evolution of a quantum state when decoherence arises from the three qubit-to-environment interactions presented before [27, 29, 74, 78]. Such channel consists of a generalized amplitude damping channel describing the thermal and relaxation interaction [74, 78] and of a pure dephasing channel [27, 29]. The action of the generalized amplitude damping channel is described by the damping parameter, $\gamma$, and the probability that the ground state is excited by finite temperature, $N$. The damping parameter relates to the relaxation time of the qubit as, $\gamma(t)=1-e^{-(2n_{\mathrm{th}}+1)t/T_{1}}$, with $t$ being the evolution time, while $n_{\mathrm{th}}$ and $N(n_{\mathrm{th}})$ depend on the temperature and the energy gap of the system [74, 78]. The Kraus operators of this channel are:
$E_{0}=\sqrt{1-N(n_{\mathrm{th}})}\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma(t)}\end{pmatrix}$, (12)
$E_{1}=\sqrt{1-N(n_{\mathrm{th}})}\begin{pmatrix}0&\sqrt{\gamma(t)}\\ 0&0\end{pmatrix}$, (13)
$E_{2}=\sqrt{N(n_{\mathrm{th}})}\begin{pmatrix}\sqrt{1-\gamma(t)}&0\\ 0&1\end{pmatrix},\text{ and}$ (14)
$E_{3}=\sqrt{N(n_{\mathrm{th}})}\begin{pmatrix}0&0\\ \sqrt{\gamma(t)}&0\end{pmatrix}$. (15)
Whenever the thermal excitation is considered negligible, $n_{\mathrm{th}}\approx 0$ and $N(n_{\mathrm{th}})=0$; and such channel reduces to an amplitude damping channel. The pure dephasing channel is described by the so called scattering probability, $\lambda$, which relates to the pure dephasing time as, $\lambda(t)=1-e^{-2t/T_{\phi}}$, with $t$ the evolution time again [29, 58]. The Kraus operators of this channel are:
$E_{0}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\lambda(t)}\end{pmatrix},\text{ and}$ (16)
$E_{1}=\begin{pmatrix}0&0\\ 0&\sqrt{\lambda(t)}\end{pmatrix}$. (17)
In this sense, the complete channel is defined by those parameters its Kraus operators are given by the serial concatenation of the damping and dephasing channels [58].